Problem 4 A) A local tree farm is trying to decide which fertilizer to use to grow its trees. It can either buy the famous Greenwood Fertilizer, which costs $5 and can grow a tree which can be sold for $8 in 2 years, or it can buy Brian and Peter’s Fertilizer which costs $4 and can grow a tree which can be sold for $10 in 5 years. (15 points)

  1. What is the NPV of each option given an annual effective interest rate of 5%? (5 points)
  2. What is the rate of return of each option? (5 points)
  3. Which option should the tree farm pick assuming an infinite time horizon? Why? (5 points)
Answer #1

Greenwood: NPV = -5+8/(1.05)^2 = 2.56
Bryan and Peter NPV = -4+10/(1.05)^5 = 3.84

Greenwood
5 = 8/(1+r)^2
r = 1.26 - 1
r = 26%
Bryan and Peter
4 = 10/(1+r)^5
r = 1.20 -1
r = 20%

Assuming infinite time horizon, farmer shall pick Greenwood as it has better returns